# AI-Generated Code Is Not Reproducible (Yet): An Empirical Study of Dependency Gaps in LLM-Based Coding Agents

### v1

@@ Introduction
Reproducibility forms the cornerstone of computational science. When researchers share code, others must be able to execute it, verify results, and build upon the work (gundersen2018on). This fundamental principle enables scientific progress, validates discoveries, and ensures reliability in production systems. Yet as we increasingly rely on Large Language Models (LLMs) to generate code, a troubling question emerges: is AI-generated code reproducible?
The computational research community already faces a reproducibility crisis. Studies show that a significant portion of published AI/ML research cannot be reproduced due to missing dependencies, incomplete environment specifications, and undocumented requirements (pineau2021improving; raff2019step; hutson2018artificial; haibe2020transparency). Now, tools like Claude Code, OpenAI Codex, and Gemini Code Assist promise to accelerate development by generating complete projects, entire applications with multiple files, configurations, and dependency specifications. But if these AI systems inherit or amplify reproducibility problems, we risk compounding the crisis rather than solving it.
@@
This reproducibility failure represents a fundamental gap in how we evaluate AI coding agents. Current benchmarks like HumanEval (chen2021evaluating) and MBPP (austin2021program) test functional correctness but assume reproducible environments already exist, a recognized bottleneck that obscures key failure modes related to environment configuration (yang2024swe; jimenez2024swebench). They ask, does this sorting algorithm work correctly? without first asking, can anyone actually run this code? This misses a critical insight: code cannot be correct if it cannot be reproduced. Reproducibility is not a secondary concern; it is a prerequisite for verification, collaboration, and deployment.
@@
We conducted a systematic reproducibility study of 300 complete projects generated by three leading state of the art LLM agents (anthropic2025claude; openai2025codex; google2025gemini). Each AI received identical prompts explicitly requesting reproducible code with complete dependency specifications. We attempted to reproduce each project in clean environments with only OS packages installed and none other than them, documenting every failure and missing requirement. This mimics the exact challenge faced by researchers trying to reproduce published code or developers attempting to use AI-generated projects. Our methodology treats reproducibility as the primary concern: can the generated code be executed by others using only what the AI provides?
Our findings reveal that AI-generated code faces a severe reproducibility crisis. Only 68.3% of projects are reproducible without manual intervention nearly one-third fail to execute as specified. The reproducibility rates vary dramatically by language: Python achieves 89.2% reproducibility, JavaScript 61.9%, and Java just 44.0%. This variation reflects how different ecosystems handle dependency specifications (decan2019empirical; kikas2017structure) and transitive requirements. Most striking is the runtime dependency explosion: projects claiming 3 dependencies actually require an average of 37 packages, and 13.5× gap between specified and actual requirements. This hidden complexity makes reproduction nearly impossible without extensive debugging.
These results demonstrate that current LLMs generate code that appears complete but lacks the specifications necessary for reproducibility. The implications extend beyond convenience. For scientific computing, non-reproducible code undermines research validity. For software development, it creates hidden technical debt (sculley2015hidden). For education, it teaches poor practices. By quantifying these reproducibility failures and identifying their patterns, we provide both a diagnosis of the current crisis and a roadmap for creating AI systems that generate truly reproducible code. The challenge is not just to generate code that works on the developer’s machine, but code that others can reliably execute, the foundation of computational reproducibility.
@@
The reproducibility crisis in AI research (gundersen2018on) now extends to the tools meant to address it. While computational research should enable perfect reproducibility (peng2011reproducible), LLM-generated code introduces new failure modes. Studies show LLMs produce inconsistent outputs (wang2025assessing) and unreliable automation (staudinger2024reproducibility). Our work reveals a more fundamental issue: even when LLMs generate complete projects with multiple files and configurations, they fail to specify dependencies needed for reproducible execution.
@@
Current benchmarks like HumanEval and MBPP (chen2021evaluating) evaluate functional correctness assuming reproducible environments exist. LiveCodeBench prevents memorization but still provides complete environments (jain2024livecodebench). These benchmarks test whether code is correct, not whether it’s reproducible. Our work evaluates complete projects not snippets asking whether they can execute at all using only what LLMs provide, making reproducibility the primary metric.
@@
Software reproducibility requires complete dependency specification. JavaScript packages average 79 transitive dependencies (decan2019empirical). While tools like ReproZip (reprozip) and SciUnit (8109156) capture execution environments for reproducibility, LLMs lack this awareness. They generate syntactically correct code but fail to specify the dependency closure required for reproduction.
@@
To understand whether AI-generated code can actually run in practice, we needed to formalize what reproducibility means in this context. We evaluated three leading LLM-based coding agents: Claude Code Agent(Opus.4.1) (anthropic2025claude), OpenAI Codex Agent(0.52.0) (openai2025codex), and Gemini Code Agent(2.5.Pro) (google2025gemini). We denote this set as $\mathcal{L}=\{L_{Claude},L_{Gemini},L_{Codex}\}$. Each agent was given the same 100 prompts, creating our standardized prompt collection $\mathcal{P}=\{p_{1},p_{2},...,p_{100}\}$.
@@ Discussion
where $n$ represents iteration count (typically 2-3 for missing dependencies). Our manual processing averaged 15 minutes per failed project, suggesting substantial aggregate time costs when scaled across development teams. For an organization deploying LLM-generated code at scale, this translates to approximately 8 hours of debugging for every 100 generated projects (31.7 projects × 15 minutes ÷ 60 = 7.9 hours). Furthermore, the cognitive load (sweller1988cognitive) of debugging AI-generated code differs from debugging human-written code. Developers must reverse-engineer the LLM’s intent without access to the original design rationale, making fixes more challenging and error-prone.

### v2

@@ Introduction
Reproducibility forms the cornerstone of computational science. When researchers share code, others must be able to execute it, verify results, and build upon the work (Gundersen et al. 2018). This fundamental principle enables scientific progress, validates discoveries, and ensures reliability in production systems. Yet as we increasingly rely on Large Language Models (LLMs) to generate code, a troubling question emerges: is AI-generated code reproducible?
The computational research community already faces a reproducibility crisis. Studies show that a significant portion of published AI/ML research cannot be reproduced due to missing dependencies, incomplete environment specifications, and undocumented requirements (Pineau et al. 2021; Raff 2019; Hutson 2018; Haibe-Kains and others 2020). Now, tools like Claude Code, OpenAI Codex, and Gemini Code Assist promise to accelerate development by generating complete projects, entire applications with multiple files, configurations, and dependency specifications. But if these AI systems inherit or amplify reproducibility problems, we risk compounding the crisis rather than solving it.
@@
This reproducibility failure represents a fundamental gap in how we evaluate AI coding agents. Current benchmarks like HumanEval (Chen et al. 2021) and MBPP (Austin and others 2021) test functional correctness but assume reproducible environments already exist, a recognized bottleneck that obscures key failure modes related to environment configuration (Yang et al. 2024; Jimenez and others 2024). They ask, does this sorting algorithm work correctly? without first asking, can anyone actually run this code? This misses a critical insight: code cannot be correct if it cannot be reproduced. Reproducibility is not a secondary concern; it is a prerequisite for verification, collaboration, and deployment.
@@
We conducted a systematic reproducibility study of 300 complete projects generated by three leading state of the art LLM agents (Anthropic 2025; OpenAI 2025; Google DeepMind 2025). Each AI received identical prompts explicitly requesting reproducible code with complete dependency specifications. We attempted to reproduce each project in clean environments with only OS packages installed and none other than them, documenting every failure and missing requirement. This mimics the exact challenge faced by researchers trying to reproduce published code or developers attempting to use AI-generated projects. Our methodology treats reproducibility as the primary concern: can the generated code be executed by others using only what the AI provides?
Our findings reveal that AI-generated code faces a severe reproducibility crisis. Only 68.3% of projects are reproducible without manual intervention nearly one-third fail to execute as specified. The reproducibility rates vary dramatically by language: Python achieves 89.2% reproducibility, JavaScript 61.9%, and Java just 44.0%. This variation reflects how different ecosystems handle dependency specifications (Decan et al. 2019; Kikas and others 2017) and transitive requirements. Most striking is the runtime dependency explosion: projects claiming 3 dependencies actually require an average of 37 packages, and 13.5× gap between specified and actual requirements. This hidden complexity makes reproduction nearly impossible without extensive debugging.
These results demonstrate that current LLMs generate code that appears complete but lacks the specifications necessary for reproducibility. The implications extend beyond convenience. For scientific computing, non-reproducible code undermines research validity. For software development, it creates hidden technical debt (Sculley et al. 2015). For education, it teaches poor practices. By quantifying these reproducibility failures and identifying their patterns, we provide both a diagnosis of the current crisis and a roadmap for creating AI systems that generate truly reproducible code. The challenge is not just to generate code that works on the developer’s machine, but code that others can reliably execute, the foundation of computational reproducibility.
@@
The reproducibility crisis in AI research (Gundersen et al. 2018) now extends to the tools meant to address it. While computational research should enable perfect reproducibility (Peng 2011), LLM-generated code introduces new failure modes. Studies show LLMs produce inconsistent outputs (Wang and Wang 2025) and unreliable automation (Staudinger et al. 2024). Our work reveals a more fundamental issue: even when LLMs generate complete projects with multiple files and configurations, they fail to specify dependencies needed for reproducible execution.
@@
Current benchmarks like HumanEval and MBPP (Chen et al. 2021) evaluate functional correctness assuming reproducible environments exist. LiveCodeBench prevents memorization but still provides complete environments (Jain et al. 2024). These benchmarks test whether code is correct, not whether it’s reproducible. Our work evaluates complete projects not snippets asking whether they can execute at all using only what LLMs provide, making reproducibility the primary metric.
@@
Software reproducibility requires complete dependency specification. JavaScript packages average 79 transitive dependencies (Decan et al. 2019). While tools like ReproZip (Chirigati et al. 2024) and SciUnit (Ton That et al. 2017) capture execution environments for reproducibility, LLMs lack this awareness. They generate syntactically correct code but fail to specify the dependency closure required for reproduction.
@@
To understand whether AI-generated code can actually run in practice, we needed to formalize what reproducibility means in this context. We evaluated three leading LLM-based coding agents: Claude Code Agent(Opus.4.1) (Anthropic 2025), OpenAI Codex Agent(0.52.0) (OpenAI 2025), and Gemini Code Agent(2.5.Pro) (Google DeepMind 2025). We denote this set as $\mathcal{L}=\{L_{Claude},L_{Gemini},L_{Codex}\}$. Each agent was given the same 100 prompts, creating our standardized prompt collection $\mathcal{P}=\{p_{1},p_{2},...,p_{100}\}$.
@@ Discussion
where $n$ represents iteration count (typically 2-3 for missing dependencies). Our manual processing averaged 15 minutes per failed project, suggesting substantial aggregate time costs when scaled across development teams. For an organization deploying LLM-generated code at scale, this translates to approximately 8 hours of debugging for every 100 generated projects (31.7 projects × 15 minutes ÷ 60 = 7.9 hours). Furthermore, the cognitive load (Sweller 1988) of debugging AI-generated code differs from debugging human-written code. Developers must reverse-engineer the LLM’s intent without access to the original design rationale, making fixes more challenging and error-prone.